HUBBLE SPACE TELESCOPE DAILY REPORT #2175 PERIOD COVERED: 0000Z (UTC) 07/28/98 - 0000Z (UTC) 07/29/98 Daily Status Report as of 210/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed FOC/96/WF/PC-2 7274 (Secular Changes in the Jet of M-87) The Faint Object Camera (f/96) and the WF/PC-2 were both used to make observations of the secular changes in the jet coming off of the active galactic nuclei of galaxy M-87. Past observations with the FOC have shown changes as small as 0.03", and provide a unique opportunity to study proper motions and structural changes of the jet. An internal flat calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.3 Completed Twelve Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. As documented in yesterday's report and HSTARs 6596 and 6598, the take data flag remained down for the sixth and seventh iterations of this proposal due to the gyro 6 motor current jump which occurred at 209/0501Z. Then, as described in HSTAR 6599 and 2.1, the acquisition/re-acquisition for the eighth through the final iteration of this proposal failed and the take data remained down. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.4 Completed Eight Sets of STIS/CCD 7910 (STIS Extragalactic Parallel Survey) The Space Telescope Imaging Spectrograph (CCD) was used to make extragalactic slitless spectrographic observations as part of the parallel high galactic latitude survey to search for emission line galaxies. Data was taken in parallel, and are not subject to proprietary time. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. As documented in yesterday's report and HSTARs 6596 and 6598, the take data flag remained down for the second iteration of this proposal due to the gyro 6 motor current jump which occurred at 209/0501Z. Otherwise, the observations were completed as planned, and no further problems were reported. 1.6 Completed STIS/CCD 7944 (STIS Sparse Field CTE Test) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a STIS Sparse Field (CTE) test. The CTE measurements are made using the "sparse field test", along both the serial and parallel axes. The test was completed as scheduled, and no anomalies were noted. 1.7 Completed Six Sets of STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. As documented in yesterday's report and HSTARs 6596 and 6598, the take data flag remained down for part of the second iteration and for all of the third iteration of this proposal due to the gyro 6 motor current jump which occurred at 209/0501Z. Then, as described in HSTAR 6599 and 2.1, the acquisition/re-acquisition for the fourth through the final iteration of this proposal failed and the take data remained down. Otherwise, the observations were completed as scheduled, and no further problems were reported. 1.8 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was to be used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. As documented in yesterday's report and HSTARs 6596 and 6598, the take data flag remained down for both iterations of this proposal due to the gyro 6 motor current jump which occurred at 209/0501Z. 1.9 Completed WF/PC-2 7386 (A Snapshot Survey of Nova Shells) The WF/PC-2 was to be used to make observational images of the super nova shell V1974-Cyg in order to detect and measure its expanding gaseous shell. These observations will trace the distribution of hot and cool gas in the nova shells, and show their degree of inhomogeneity on both large and small scales. As documented in yesterday's report and HSTARs 6596 and 6598, the take data flag remained down this proposal due to the gyro 6 motor current jump which occurred at 209/0501Z. 1.10 Completed NIC/1/2/STIS/CCD/MA1 7565 (Comparison of Protoplanetary Disks) The Near Infrared Camera (Cameras One and Two) and the Space Telescope Imaging Spectrograph (CCD and MA1) were to be used to make observations of the protoplanetary system around star HD-163296. This is to search for planets and disk structures around nearby bright stars, comparing an example of a solar type and a hotter star, each with a main sequence and a pre-main sequence star, using broad band coronagraphic imaging using the STIS CCD. Multiple roll angles were to be used to minimize systematic errors and image artifacts. As documented in yesterday's reporta and HSTARs 6596 and 6598, the take data flag remained down for this proposal due to the gyro 6 motor current jump which occurred at 209/0501Z. 1.11 Completed Five Sets of STIS/CCD 7969 (STIS CCD G750L Fringe Flats at Intermediate MSM Positions) The Space Telescope Imaging Spectrograph (CCD) was used to investigate the flat-field fringing in the G750L grating at a number of wavelength settings. WAVECALs were also taken at the same gain and wavelength setting. As documented in yesterday's report and HSTARs 6596 and 6598, the take data flag remained down for most of the first iteration of this proposal due to the gyro 6 motor current jump which occurred at 209/0501Z. Then, as described in HSTAR 6599 and 2.1, the acquisition/re-acquisition for the fourth iteration of this proposal failed and the take data remained down. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.12 Completed STIS/CCD/MA1/WF/PC-2/NIC/1/2 7570 (Spectra of Moderate Z Cluster Galaxies) The Space Telescope Imaging Spectrograph (CCD and MA1), the WF/PC-2 and the Near Infrared Camera (Cameras One and Two) were all used to make spectrographic observations of the galaxy A2390-3002. The observations will be used to study the morphology and spectrographic gradients in the cD class galaxy and to look for the cooling flow in O-VI line data. The observations were completed as planned, and no anomalies were noted. 1.13 Completed NIC/1/2/3 7896 (A NIR Snapshot Survey of Ultraluminous IR Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make snapshot survey observations of the ultraluminous infrared galaxy CL-ULIR222. Nearly all of these galaxies show evidence for a recent tidal interaction, and we will search for evidence of the merger progenitors near the center of each galaxy, a task made significantly easier with NIR imaging. The observations were completed as planned, and no anomalies were noted. 1.14 Completed FOC/96 6713 (Geometric Measurement of Galaxy Distance) The Faint Object Camera (f/96) was to be used to make observations of supernova remnant SN1970G in order to make a distance calibration for the galaxy in which it is located. The supernova is being used as a powerful new geometric method for the determination of a galaxy's distance, by obtaining high spatial resolution imaging polarization of the candidate supernova light echoes. As described in HSTAR 6599 and 2.1, the acquisition/re-acquisition for this proposal failed and the take data remained down. 1.15 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was to be used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. As described in HSTAR 6599 and 2.1, the acquisition/re-acquisition for this proposal failed and the take data remained down. 1.16 Completed WF/PC-2 8059 (POMS Test Proposal: WF-2 Targeted Parallel Archive Proposal) The WF/PC-2 was to be used to make observations in parallel with the prime science observations. The parallel opportunities available with WF/PC-2 in the neighborhood of bright galaxies are treated in a slightly different way from the normal pure parallels. Local Group galaxies offer the opportunity for a closer look at young stellar populations. As described in HSTAR 6599 and 2.1, the acquisition/re-acquisition for this proposal failed and the take data remained down. 1.17 Completed STIS/CCD 7867 (NICMOS Imaging of the Dusty CfA Seyfert Nuclei) The Space Telescope Imaging Spectrograph (CCS) was used to make observations of the Seyfert Galaxy MRK334. These data will allow us to determine if the Seyfert Type 1/2 dichotomy is best explained by the unified model, the evolutionary picture, or a combination of both of these scenarios. The observations were executed as scheduled, and no anomalies were reported. 1.18 Completed Two Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.19 Completed NIC/2 6767 (A Search for Brown Dwarfs and Luminous Young Planets in the Hyades) The Near Infrared Camera (Camera Two) was used to make observations of the brown dwarf star RHY083 which is located in the Hyades Cluster. This is to search for substellar companions orbiting low-mass Hyades M class dwarf stars. Differential imaging was used using methane and water band (on/off) filter pairs. The observations were completed as planned, and no problems were reported. 1.20 Completed WF/PC-2 7427 (Saturn's Rings) The WF/PC-2 was used to make imaging observations of both the east and west apsides of Saturn's rings. Saturn's rings are a unique dynamical laboratory and continue to present challenging problems which require additional data to attack. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 12 Successful : 9 Scheduled Reacquisitions : 5 Successful : 4 Per HSTAR 6599, there were FHST update failures at 209/143349Z and at 209/143634Z followed by an acquisition failure at 209/1441Z and a reacquisition failure at 209/1618Z. All were caused by a V3 attitude error. Proposals described in 1.3, 1.7, 1.11, 1.14, 1.15, and 1.16 were affected. HSTAR 6601 describes a guide star acquisition failure at 210/031023Z when the search radius limit was exceeded. The proposals (all occurring during the period of the next report) affected were: STIS 7969 ((STIS CCD G750L Fringe Flats at Intermediate MSM Positions), STIS 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup), NIC 7907 (NICMOS Pure Parallel Winter 97-98), WF/PC-2 6713 (Geometric Measurement of Galaxy Distance), and WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal). HSTAR 6602 describes another guide star acquisition failure at 210/051924Z when the search radius limit was exceeded due to a large attitude error. The proposals (occurring during the period of the next report) affected were: STIS 7439 (H-Lyman Alpha Emission from the Upper Atmosphere of Uranus) and NIC 7896 (A NIR Snapshot Survey of Ultraluminous IR Galaxies). 2.2 FHST Updates: Scheduled : 19 Successful : 17 2.3 Operations Notes: While attempting to deal with the attitude errors encountered for this operational day, three high/low gyro bias updates and two attitude reference updates were performed. Using ROP SR-1, the SSR EDAC error counter was cleared. FOC errors were reset at 209/1744Z per ROP NS-6. As documented in HSTAR 6600, the STIS MCE-1 reset at 209/170448Z while operating in low voltage. Using ROP NS-12, the flight software errors were cleared at 209/1745Z. The STIS MAMA-1 was recovered via the SMS at 210/0554Z. ************************************************************************* PSEA was reconfigured for RGAs 3, 1 ,5, and 2 at 209/2323Z per an operations request. ************************************************************************* 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH